Mzomuhle takes out a personal loan of R439200,00 to help finance the building of his holiday house.The terms of his loan specify equal three-monthly repayments over five years, with13,2% interest per annum, compounded quarterly.The first payment is made three months after the loan was taken out. First find the size of the three-monthly payments.Then considering the amortisation schedule,find the total interest charged over the first year of repayments.


1
Expert's answer
2021-05-24T18:53:03-0400

We will find it by the formula:


"А=S\times(\dfrac{r}{(1+r)^n-1})"


S= 439 200


Since the payment is quarterly, we will change the rate and the number of periods accordingly


"r=\dfrac{0.132}{4}=0.033"


"n=4\times5=20"


"А=S\times(\dfrac{r}{(1+r)^n-1})=439 200\times(\dfrac{0.033}{(1+0.033)^{20}-1})=15 852.40"